# Visibility into AI Agents

### v4

@@ Abstract.
copyright: noneconference: 2024 ACM Conference on Fairness, Accountability, and Transparency; June 3–6, 2024; Rio de Janeiro, Brazilbooktitle: 2024 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’24), June 3–6, 2024, Rio de Janeiro, Brazilcopyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: FAccT 2024; June 03–06, 2024; Rio de Janeiro, Brazilisbn: 978-1-4503-XXXX-X/18/06
@@ 1.1. Risks of AI Agents
Rather than provide an exhaustive taxonomy of risks from AI agents, [^3] See Critch and Russell (2023); Weidinger et al. (2023); Shelby et al. (2023) for taxonomies of risks from AI systems and (Chan et al., 2023b; Shavit et al., 2023) for further discussion of risks from AI agents. we highlight certain agent-specific risks. In comparison to risks from other AI systems, these risks focus on the potential for increasingly agentic systems to substitute for, rather than complement, human labour (Korinek and Juelfs, 2022) and remove humans from the loop (Chan et al., 2023b). Without a human in the loop, agents may take multiple consequential actions in rapid succession and bring about significant impacts before a human notices. The ability to remove humans from the loop also means that an agent’s task performance is less limited by the expertise of its user, compared to a situation where user must guide an AI system’s actions or take actions themself.
@@ 1.1.5. Sub-Agents
Agents could instantiate more agents to achieve user goals, which may magnify several of the risks discussed so far. It may be advantageous for an agent to create potentially specialized and more efficient sub-agents, especially if doing so is cheap and fast. For example, an agent could call copies of itself through an API, or itself train, fine-tune, or otherwise program another agent to accomplish (components of) a task. Sub-agents could be problematic because they introduce additional points of failure; each sub-agent may itself malfunction, be vulnerable to attack, or otherwise operate in a way contrary to the user’s intentions. To stop an agent from causing harm, one would have to intervene not only on the agent, but also on any relevant sub-agents (Carey and Everitt, 2023; Soares et al., 2015). Yet, detecting harmful behaviour and intervening may be difficult because we lack methods for determining when an agent has created a sub-agent. Additionally, we would need to ensure that safety constraints imposed on the original agent faithfully transfer to sub-agents.
@@ 2. Definitions
Agency is the degree to which an AI system acts directly in the world to achieve long-horizon goals, with little human intervention or specification of how to do so. An (AI) agent is a system with a relatively high degree of agency; we consider systems that mainly predict without acting in the world, such as image classifiers, to have relatively low degrees of agency. Examples of agents could include reinforcement learning systems (Team et al., 2023a; Reed et al., 2022) that interact extensively with the real world [^6] Including the physical environment but also digital environments such as online platforms. or more capable versions of language models with tool access (OpenAI, 2023a; Schick et al., 2023; Ruan et al., 2023; Bran et al., 2023). We do not consider existing foundation models themselves to be agents. Our definitions compress the characterization of agency in Chan et al. (2023b), which points to four axes: the degree to which the system’s behaviour is specified, the degree to which the system’s behaviour is goal-directed, the degree to which the system has a direct impact in the world, and the degree to which the system can achieve goals over long time-horizons. For the purposes of this paper, we use agent and agentic system interchangeably.
@@
The (agent) deployer is the entity that operates an AI system and serves it to users. The deployer may not be the same as the developer(s). For example, Microsoft deploys OpenAI’s systems into its products (Warren, 2024), but did not develop GPT-4. A deployer may provide access to an agent in one of two ways. First, the deployer may serve a foundation model which users may combine with other components to make a moreThe system may not be completely autonomous since user approval may still be required for certain actions. agentic system. For example, users may use the scaffolding framework AutoGPT (Richards, 2023) to chain calls to GPT-4 (OpenAI, 2023b) and provide the model with tool access. Indeed, popular scaffolding frameworks depend upon an underlying foundation model (Wu et al., 2023; Richards, 2023). Second, the deployer may provide an agent or may furnish ways for users to make a provided system more agentic. For instance, OpenAI lets users build and share custom agents (OpenAI, 2024, 2023c) augmented with a variety of tools, including browsing, using Google Drive apps, and coding (OpenAI, 2023a).
@@ 3.1. Agent Identifiers
The ability to identify agents could be useful to several actors. A regulator could require AI agents to identify themselves as non-humans during interaction with humans (Frazier, 2023; Baker, 2023), similar to bot disclosure laws (DiResta, 2019). [^8] Note that not all non-human activities may come from AI agents. For example, consider currently automated trading activity or ads auctions. Members of the general public may wish to know whether they have interacted with AI agents. Summary statistics about when agent identifiers are visible could inform governments and the general public about the extent to which agents operate in high-risk settings (noa, 2021). Identifiers for when agents send requests to tools or services providers could help to identify significant actions, such as when agents transfer sensitive information. A service provider may even reject a request absent certain guarantees attached to the identifier, such as those related to the security of the agent. Unique identifiers for each AI agent could facilitate accountability by linking an action to an AI agent and its user, developer(s), and deployer.
@@ Acknowledgements.
We are extremely grateful to the following people for insightful feedback and conversations over the course of writing this work: Charlotte Siegmann, Chinmay Deshpande, Nathan Barnard, Leonie Koessler, Kevin Frazier, Matthijs Maas, Micah Carroll, Yonadav Shavit, Merlin Stein, Ben Bucknall, Shalaleh Rismani, Paul Crowley, Gabriel Recchia, Janet Egan.

### v5

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: The 2024 ACM Conference on Fairness, Accountability, and Transparency; June 3–6, 2024; Rio de Janeiro, Brazilbooktitle: The 2024 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’24), June 3–6, 2024, Rio de Janeiro, Brazildoi: 10.1145/3630106.3658948isbn: 979-8-4007-0450-5/24/06ccs: Computing methodologies Artificial intelligenceccs: Applied computing Law
@@ 1.1. Risks of AI Agents
Rather than provide an exhaustive taxonomy of risks from AI agents, [^3] See Critch and Russell (2023); Weidinger et al. (2023); Shelby et al. (2023) for taxonomies of risks from AI systems and (Chan et al., 2023b; Shavit et al., 2023) for further discussion of risks from AI agents. we highlight certain agent-specific risks. In comparison to risks from other AI systems, these risks focus on the potential for agents to remove humans from the loop (Korinek and Juelfs, 2022; Chan et al., 2023b). Without a human in the loop, agents may take multiple consequential actions in rapid succession and bring about significant impacts before a human notices. The ability to remove humans from the loop also means that an agent’s task performance is less limited by the expertise of its user, compared to a situation where user must guide an AI system’s actions or take actions themself.
@@ 1.1.5. Sub-Agents
Agents could instantiate more agents to accomplish (components of) a task, which may magnify several of the risks discussed so far. It may be advantageous for an agent to create potentially specialized and more efficient sub-agents, especially if doing so is cheap and fast. For example, an agent could call copies of itself through an API, or itself train, fine-tune, or otherwise program another agent. Sub-agents could be problematic because they introduce additional points of failure; each sub-agent may itself malfunction, be vulnerable to attack, or otherwise operate in a way contrary to the user’s intentions. Stopping an agent from causing further harm might involve intervening not only on the agent, but also on any relevant sub-agents (Carey and Everitt, 2023; Soares et al., 2015). Yet, this process may be difficult because we lack methods for determining when an agent has created a sub-agent.
@@ 2. Definitions
Agency is the degree to which an AI system acts directly in the world to achieve long-horizon goals, with little human intervention or specification of how to do so. An (AI) agent is a system with a relatively high degree of agency; we consider systems that mainly predict without acting in the world, such as image classifiers, to have relatively low degrees of agency. Examples of agents include reinforcement learning systems (Team et al., 2023a; Reed et al., 2022) that interact extensively with the real world [^6] Including the physical environment but also digital environments such as online platforms. or more capable versions of language models with tool access (OpenAI, 2023a; Schick et al., 2023; Ruan et al., 2023; Bran et al., 2023). We do not consider existing foundation models themselves to be agents. Our definitions compress the characterization of agency in Chan et al. (2023b), which points to four axes: the degree to which the system’s behaviour is specified, the degree to which the system’s behaviour is goal-directed, the degree to which the system has a direct impact in the world, and the degree to which the system can achieve goals over long time-horizons. For the purposes of this paper, we use agent and agentic system interchangeably.
@@
The (agent) deployer is the entity that operates an AI system and serves it to users. The deployer may not be the same as the developer(s). For example, Microsoft deploys OpenAI’s systems into its products (Warren, 2024), but did not develop GPT-4. A deployer may provide access to an agent in one of two ways. First, the deployer may serve a foundation model which users may combine with other components to make a more [^7] The system may not be completely autonomous since user approval may still be required for certain actions. agentic system. For example, users may use the scaffolding framework AutoGPT (Richards, 2023) to chain calls to GPT-4 (OpenAI, 2023b) and provide the model with tool access. Indeed, popular scaffolding frameworks depend upon an underlying foundation model (Wu et al., 2023; Richards, 2023). Second, the deployer may provide an agent or may furnish ways for users to make a provided system more agentic. For instance, OpenAI lets users build and share custom agents (OpenAI, 2024, 2023c) augmented with a variety of tools, including browsing, using Google Drive apps, and coding (OpenAI, 2023a).
@@ 3.1. Agent Identifiers
The ability to identify agents could be useful to several actors. A regulator could require AI agents to identify themselves as non-humans during interaction with humans (Frazier, 2023; Baker, 2023), similar to bot disclosure laws (DiResta, 2019). [^9] Note that not all non-human activities may come from AI agents. For example, consider currently automated trading activity or ads auctions. Members of the general public may wish to know whether they have interacted with AI agents. Summary statistics based on agent identifiers could inform governments and the general public about the extent to which agents operate in high-risk settings (noa, 2021). Identifiers for when agents send requests to tools or services providers could help to identify significant actions, such as when agents transfer sensitive information. A service provider may even reject a request absent certain guarantees attached to the identifier, such as those related to the security of the agent. Unique identifiers for each AI agent could facilitate accountability by linking an action to an AI agent and its user, developer(s), and deployer.
@@ Acknowledgements.
We are grateful to the following people for insightful feedback and conversations over the course of writing this work: Charlotte Siegmann, Chinmay Deshpande, Nathan Barnard, Leonie Koessler, Kevin Frazier, Matthijs Maas, Micah Carroll, Yonadav Shavit, Merlin Stein, Ben Bucknall, Shalaleh Rismani, Paul Crowley, Gabriel Recchia, Janet Egan.